OpenStudy (anonymous):

How to find the integral of 4x/(x^2-1)^2 dx This is what I did but I'm not sure about it

OpenStudy (anonymous):

4 integral x(x^2-1)^-2 dx

OpenStudy (anonymous):

U=x^2-1 Du=2x dx

OpenStudy (anonymous):

just let u = x^2 -1, you have du =2x dx, that's enough

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

And what do I do next?

OpenStudy (anonymous):

du = 2x dx --> xdx =1/2 du, ok?

OpenStudy (anonymous):

So it's 4 times 1/2 from the outside

OpenStudy (anonymous):

you can simplify it , right?